OpenStudy (anonymous):

Identify the center of the circle. (x+3)^2+(y-2)^2=1

OpenStudy (angela210793):

(-3:2)

OpenStudy (angela210793):

the eq of the circle whose center isn't in the origin is (x-x1)^2+(y-y1)^2=r^2 so x1=-3 and y1=2

OpenStudy (anonymous):

THANKS SOOOO MUCH

OpenStudy (angela210793):

yw :)